Calculate the rate of heat conduction out of the human body, assuming that the core internal temperature is 31.97 ºC, the skin temperature is 20.96 ºC, the thickness of the tissues between averages 1.51 cm, and the surface area is 1.73 m2. (skin = 0.33W/m⋅°C)


1
Expert's answer
2021-03-29T09:06:53-0400
"P=\frac{kA(T_h-T_c)}{d}\\
P=\frac{(0.33)(1.73)(31.97-20.96)}{0.0151}=416\ W"